I am a number between 500 and 600. my ones digit is 5. my tens digit is the difference between my ones and hundreds digit. Who am I?
Answer by JBarnum(2146)

I am a number between 500 and 600. 500 my ones digit is 5.(xy5)
my tens digit is the difference between my ones and hundreds digit.
ones digit=z tens digit=y hundreds digit=x
y=z-x x=5 and z=5
y=5-5
y=0
so the number xyz is 505